40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1281     Entry time: Fri Feb 6 16:20:52 2009
Author: Yoichi 
Type: Update 
Category: PSL 
Subject: MOPA current slider fixed 

I fixed the broken slider to change the current of the PA.

The problem was that the EPICS database assigned a wrong channel of the DAC to the slider.

I found that the PA current adjustment signal lines are connected to the CH3 &CH4 of VMIC4116 #1. However in the database file (/cvs/cds/caltech/target/c1psl/psl.db), the slider channel (C1:PSL-126MOPA_DCAMP) was assigned to CH2. I fixed the database file and rebooted c1psl. Then the PA current started to follow the slider value.

I moved the slider back and forth by +/-0.3V while the ISS loop was on. I observed that the amount of the low frequency fluctuation of the MOPA power changed with the slider position. At some current levels, the ISS instability problem went away.

Kakeru is now taking open-loop TFs and current shunt responses at different slider settings.

ELOG V3.1.3-